# APEX-MR: Multi-Robot Asynchronous Planning and Execution for Cooperative Assembly

### v1

@@ IV APEX-MR: Asynchronous Planning and Execution for Multi-Robot System
In this section, we introduce APEX-MR, a framework for asynchronously coordinating the task plan, motion, and execution of a multi-robot system to accomplish cooperative assembly tasks. Fig. 2 provides an overview of the three stages of APEX-MR. The first task planning stage is discussed in Sec. IV-A. Given the input discussed in Sec. III, a sequential task plan and the corresponding task graph, as shown in Fig. 2 (a), are generated from the assembly using ILP. In the next stage explained in Sec. IV-B, the motion plan for each task is generated sequentially with a single-robot motion planner (see Fig. 2 (b. Sec. IV-C presents the last and most crucial stage, which converts the sequential plan to a multi-modal TPG (e.g. Fig. 2 (c)) that can be executed safely, asynchronously, and efficiently in real robots.
For notation, we use $i\in[1,N]$ to index over robots, $j\in[1,N_{a}]$ to index over assembly steps in $A$, $m$ to index over tasks for each robot, and $k\in[1,N_{b}]$ to index over usable objects in $\mathcal{B}$.
@@ IV-A Task Planning
Given the assembly sequence $A$, task planning aims to construct a sequential task plan that includes robot assignment, robot target pose assignment, and object assignment. Specifically, for each step $a_{j}$ in the assembly sequence, the task planner must assign one responsible robot, a support robot if required, and an object $b_{k}$ of the correct type. In addition, the task planner must select the feasible grasp pose, or a composite motion of multiple poses, to perform necessary manipulation skills for each $a_{j}$. Fig. 3 provides an overview of all the decisions made. These decisions directly affect the feasibility of motion planning, as task planning determines whether there are collision-free and deadlock-free paths.
@@
Task Definition We define a task $\mathcal{T}^{i}$ as a piece of work that requires robot $i$ to either transit to a goal pose or to perform some manipulation skill $s$ to achieve a goal constraint. For example, a task may involve a robot arm picking up an object, supporting a structure, receiving an object handed from another robot, retracting to its HOME pose, etc. In the context of LEGO assembly, each assembly task $a_{j}$ is broken down into a sequence of tasks. As shown in Fig. 1 (a) and (b), a robot must first transit to the initial location of an object, pick the object, transit the object to the location for assembly, place down the object on the target structure, and finally transit back to the robot’s HOME. For an assembly step that requires collaborative assembly, the support robot is also assigned a sequence of tasks, such as transiting to the structure and supporting it in Fig. 1 (c) and (e). If a reorientation is required as illustrated in Fig. 1 (d), the support robot will be assigned a transit and a pick task to collect the object, and a handover task to the other robot so that the object can be placed up to the structure.
@@
Task Graph A task graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a direct acyclic graph that represents the ordered set of tasks for all robots to complete the final assembly. The task graph also implicitly represents the movement of manipulated objects, the evolving environment and collision scene, as well as the kinematic switches of the robots. A node is either a task node or an object node. A task node represents a task $\mathcal{T}^{i}_{m}\in\mathcal{V}$ for robot $i$, where $m$ is the index. Each task node also contains a robot target pose for this task. An object node represents an object $b_{k}$ and its poses. An edge from one task node to another in the graph, $\mathcal{T}^{i}_{m}\rightarrow\mathcal{T}^{i^{\prime}}_{m^{\prime}}$, represents a precedence constraint, that states a dependency where $\mathcal{T}^{i}_{m}$ must finish before $\mathcal{T}^{i^{\prime}}_{m^{\prime}}$ can start. An edge from an object node to a task node, $b_{k}\rightarrow\mathcal{T}^{i}_{m}$, means that the object $b_{k}$ is kinematically attached to the robot $i$ at the beginning of this task. In contrast, an edge from a task node to an object node, $\mathcal{T}^{i}_{m}\rightarrow b_{k}$, represents that an object would be detached from robot $i$ after this task ends. Each robot $i$ has $M^{i}$ tasks. A task graph itself does not limit if the tasks must be executed sequentially, synchronously, or asynchronously.
@@
Approach The main idea of our approach is to find a turn-based, sequential task plan according to the assembly sequence $A$, which is itself sequential. Only one robot is actively moving or executing skills at any time, while the other robot waits. Each robot returns to its HOME pose at the end of completing an assembly step.
@@
Algorithmically, an ILP jointly optimizes robot assignment, object assignment, and target robot poses. A set of binary decision variables assigns a robot $i$ to an assembly step $a_{j}$ using the object $b_{k}$ and a feasible grasp pose. Another set of binary decision variables denotes the assignment of support robot and support poses. We precompute feasible robot poses for grasping all available objects at initial and assembled positions, and support poses if necessary. The cost for assigning a robot $i$, an object $b_{k}$, and a corresponding grasp pose to an assembly step $a_{j}$ is estimated by the sum of transit distances necessary for this assembly step. The ILP finds the best set of assignments that minimizes the sum of costs to complete the assembly and an auxiliary term for load balancing while ensuring the object type and support robot requirements are met. More details of the ILP formulation are discussed in Sec. -A.
Combined with the precomputed robot poses, the optimized assignment gives a complete set of robot, object, and grasp pose assignment in each step. We then construct a corresponding sequential task plan and task graph (e.g., Fig. 2 (a)) with all inter-robot task dependencies and object relationships. A sequence of tasks for the assigned robot is added for each assembly step, and two object nodes are added to the task graph and connected to the pick task and place task nodes to indicate their attachment and placement, respectively. If a collaborative assembly is required, a sequence of support tasks is added to the support robot to be completed first. Inter-robot task dependencies are added to the task graph to constrain that the support task must precede any place task, and the following task after support can only start after the place task finishes. If two consecutive assembly steps $a_{j}$ and $a_{j+1}$ are assigned to different robots, a task dependency is added to ensure the place task for $a_{j+1}$ must wait for the place task for $a_{j}$.
Compared to other MR-TAMP and assembly methods such as [14, 18, 42, 5], APEX-MR generates a sequential plan first. This has two advantages: (1) It is easy to reason about inter-robot collision for scheduling tasks and avoids expensive feasibility checks needed for parallel task execution; (2) The complexity of motion planning is significantly reduced since each task becomes a single-robot planning problem, which avoids solving a challenging and time-consuming multi-robot arm motion planning problem. It is worth noting that, while the sequential plan might seem inefficient for a multi-robot system, our TPG execution framework will post-process this plan to enable efficient parallel execution.
@@ IV-B Motion Planning
Once the sequential task plan for each robot is determined, motion planning becomes straightforward. As illustrated in Fig. 2 (b), APEX-MR iterates over each task and use a single-robot RRT-Connect algorithm to plan the trajectory for this task when other robots are waiting. This is feasible because all other robots would be at a nonblocking stationary pose, i.e., HOME. A reference trajectory is also generated for tasks executed by specific manipulation skills based on the grasp/support pose.
The motion planner generates a planned trajectory for every robot and every task. Each trajectory $\tau^{i}_{m}$ for robot $i$ and task $\mathcal{T}_{m}^{i}$ is represented with a sequence of uniformly timestamped poses, $\{(C_{n}^{i},t^{i}_{n})\},n\in[N^{i}_{start,m},N^{i}_{end,m}]$ with step size $\Delta t$. $N^{i}_{start,m}$ and $N^{i}_{end,m}$ are the first and last index of the trajectory $\tau^{i}_{m}$, $N^{i}_{start,1}=1$, and $N^{i}_{start,m+1}=N^{i}_{end,m}+1$ for $m=1,\dots M-1$. For each task, the planner first generates a sequence of poses ${C^{i}_{n}}$ for the robot $i$, then determines corresponding timesteps. To ensure that each robot takes turns to complete its task according to the task sequences, the initial timestep of each task is equal to the last timestep of the previous task in the sequential plan. Then, the timestep $t_{n}^{i}$ for each pose of the same task is set based on the maximum robot velocity, i.e., $t^{i}_{n}=t^{i}_{n-1}+d(C^{i}_{n},C^{i}_{n-1})/v_{max}$. Since RRT-Connect may produce jerky and long trajectory, we use a randomized shortcutting algorithm to smooth suboptimal trajectories. The output of motion planning should be a sequence of trajectories for each task, i.e., $\tau^{i}_{m},\;\;\forall i\in N,\;\forall m\in M^{i}$.
@@ IV-C Asynchronous Execution
TPG Definition APEX-MR uses a multi-modal temporal plan graph (TPG) to represent an execution schedule for the team of robot manipulators. As illustrated in Fig. 2 (c), multi-modal TPG is a directed acyclic graph $G=(V,E)$ with two types of node. A pose node $v_{n}^{i}$ corresponds to a configuration $C_{n}^{i}$ on the trajectory of a transit task for robot $i$. A skill node $\tilde{v}^{i}_{n}$ represents a manipulation skill (e.g., pick, handover, support, etc) that will be executed by some robot controller or policy. The skill node also contains a reference trajectory generated in motion planning. A type-1 edge $(v_{n}^{i}\rightarrow v_{n+1}^{i})$ connects two nodes from the same robot $i$ and indicates the order between two nodes. A type-2 edge $(v_{n}^{i}\rightarrow v_{n^{\prime}}^{i^{\prime}})$ represents an inter-robot precedence order that constrains the robot $i^{\prime}$ to wait for robot $i$ to reach $v_{n}^{i}$ (i.e., reach the pose $C_{n}^{i}$ or complete the manipulation skill) before moving to $v_{n^{\prime}}^{i^{\prime}}$. A type-2 edge can be added for both task dependencies and motion dependencies. In contrast to the TPG defined in [20], a multi-modal TPG combines a TPG with a task graph and assigns a corresponding task $\mathcal{T}_{m}^{i}$ to each node $v_{n}^{i}$. Since there are kinematic switches and changes to the collision environment, each node also contains the robot kinematic (i.e., attached object), which will be important for the TPG construction process.
@@
Input: A task graph $\mathcal{G}$and all robot trajectories $\tau^{i}_{m}$robot $i=1,\dots,N$task $\mathcal{T}_{m}^{i}=\mathcal{T}^{i}_{1},\dots,\mathcal{T}_{M^{i}}^{i}$task $\mathcal{T}_{m}^{i}$is a transit to a goal pose
1:Add a sequence of nodes $\{v^{i}_{n}\}$ from $\tau^{i}$
@@
9:Simplify TPG with transitive reduction.
@@
Building a multi-modal TPG The process of constructing a TPG from a multi-robot motion plan can be interpreted as converting from the sequential robot trajectories to temporally dependent robot schedules. On a high level, inter-robot task dependencies are copied as type-2 edges for TPG, and type-2 edges for motion dependencies are constructed by scanning for collisions between all pair of TPG nodes. A key benefit of this partial-order representation is that TPG does not specify a fixed time between two consecutive nodes, and thus allows execution delays. We outline the detailed procedure below.
@@
The construction process begins by creating the nodes and type-1 edges. For each transit task $\mathcal{T}_{m}^{i}$, a pose node $v_{n}^{i}$ is constructed for each configuration $C_{n}^{i}$ of the trajectory $\tau_{m}^{i}$. A skill node $\tilde{v}_{n}^{i}$ is created for every manipulation task. The timestamp in the input trajectory $t_{n}^{i}$ for each node $v_{n}^{i}$ is also recorded. Each node $v_{n}^{i}$ is then connected to its successor $v_{n+1}^{i}$ by a type-1 edge and the end node of a task $v_{end,m}^{i}$ is connected to the start node of the next task $v_{start,m+1}^{i}$. When a robot waits at a node $v_{m}^{i}$, that is $C_{m}^{i}=C_{m-1}^{i}$, the node $v_{m}^{i}$ is removed because removing wait action does not change the robot path or any temporal dependency in the TPG.
@@
Next, all the type-2 edges are identified the TPG. First, all inter-robot task dependencies $(\mathcal{T}_{m}^{i}\rightarrow\mathcal{T}_{m^{\prime}}^{i^{\prime}})\in\mathcal{E}$ from the task graph $\mathcal{G}$ are added as type-2 edges to the TPG. Specifically, for every edge in the task graph $\mathcal{T}_{m}^{i}\rightarrow\mathcal{T}_{m^{\prime}}^{i^{\prime}}$, a type-2 edge is added from the beginning of task $\mathcal{T}_{m+1}^{i}$, $v_{{start,m+1}}^{i}$, to the beginning of task $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$, $v_{{start,{m^{\prime}}}}^{i^{\prime}}$. This type-2 edge ensures that robot $i$ only starts $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$ after robot $i^{\prime}$ finishes task $\mathcal{T}_{m}^{i}$.
@@
Then, motion-level dependencies are identified by iterating over all pairs of nodes in the TPG with a double loop iterating over $(i,i^{\prime})$. Robot kinematics can be incrementally changed if node $v_{n}^{i}$ is the start of a new task that attaches or detaches an object. For each node $v_{n}^{i}$, it is checked against all other nodes $v_{n^{\prime}}^{i^{\prime}}$ from a different robot ($i\neq i^{\prime}$) and has an earlier timestamp $t^{i^{\prime}}_{n^{\prime}}\leq t_{n}^{i}$. Collision checking for pose nodes means that the robot links and attached objects at the corresponding poses $C_{n}^{i}$ and $C_{n^{\prime}}^{i^{\prime}}$ are checked. For skill nodes, all robot poses on the reference trajectory have to be collision-free simultaneously. If there is any collision between $v_{n^{\prime}}^{i^{\prime}}$ and $v_{n}^{i}$, a type-2 edge is added from the earlier node’s successor, $v_{n^{\prime}+1}^{i^{\prime}}$, to the later node, $v_{n}^{i}$. This way, if $v_{n^{\prime}}^{i^{\prime}}$ has a smaller timestamp than $v_{n}^{i}$ in the input sequential plan, i.e., $t_{n^{\prime}}^{i^{\prime}}<t_{n}^{i}$, $v_{n^{\prime}}^{i^{\prime}}$ must still be executed before $v_{n}^{i}$ can start, avoiding any potential collisions. When iterating $n^{\prime}$, it is unnecessary to check collisions for any node $v_{n^{\prime}}^{i^{\prime}}$ that has a larger timestamp than $v_{n}^{i}$, i.e., $t_{n^{\prime}}^{i^{\prime}}\geq t_{n}^{i}$, because those would be checked when the iterated robot pair ($i$, $i^{\prime}$) are swapped. Also, if the current $v_{n^{\prime}}^{i^{\prime}}$ is a predecessor of $v_{n}^{i}$ in the graph, $v_{n}^{i}$ already waits for $v_{n^{\prime}}^{i^{\prime}}$ and avoid collisions, so it becomes unnecessary to check them again.
@@
Figure 4: Illustration of the execution setup. TPG maintains and controls the execution schedule of all robots on a central server, gradually sending new trajectory segments that can be safely executed. Each robot maintains a controller-sensing loop independently while updating its state with the central server.
Further Optimization An optional step to further reduce the execution makespan and smooth the trajectory is to skip the intermediate transition to HOME after every assembly step. We use the following shortcutting algorithm, similar to the strategy implemented in [41], to achieve that while maintaining a collision- and deadlock-free plan.
The anytime algorithm works by randomly sampling two nodes of the TPG ($v_{n}^{i},v_{n^{\prime}}^{i}$) from the same transit task $\mathcal{T}_{m}^{i}$, checking whether connecting them in a shortcut is feasible. Consecutive transit tasks passing through the robot’s HOME pose are merged as a single task. This allows the HOME pose to be skipped. A shortcut path directly interpolates between $C_{n}^{i}$ and $C_{n^{\prime}}^{i}$ to generate a sequence of poses with the same step size $\Delta t$. The shortcut must be collision-checked against any independent nodes (i.e., nodes that are not predecessors of $v_{n}^{i}$ or successors to $v_{n^{\prime}}^{i}$ in the TPG). On a multi-modal TPG, the collision checking must include any attached objects to the robot, as well as any independent object nodes on a task graph (i.e., object nodes that are not predecessors or successors of the current task). Once a valid shortcut is found, the original nodes between $v_{n}^{i},v_{n^{\prime}}^{i}$ are replaced with new nodes corresponding to the shortcut. If adding a valid shortcut from $v_{n}^{i}$ to $v_{n^{\prime}}^{i}$ skips any outgoing edges between these two nodes, the start node of these outgoing edges is moved to $v_{n}^{i}$. If any incoming edges are skipped, then the end node of these incoming edges is moved to $v_{n^{\prime}}^{i}$. These two steps ensure that any dependencies before adding a shortcut still exist after, and the TPG remains collision-free. Since each shortcut is collision-free, no new dependencies in the TPG are introduced and TPG remains deadlock-free. The shortcutting algorithm keeps identifying valid shortcuts until a user-defined time limit is reached, removing redundant transitions to HOME poses in the process.

### v2

@@ IV APEX-MR: Asynchronous Planning and Execution for Multi-Robot System
In this section, we introduce APEX-MR, a framework for asynchronously coordinating the task plan, motion, and execution of a multi-robot system to accomplish cooperative assembly tasks. Fig. 2 provides an overview of the three stages of APEX-MR. The first task planning stage is discussed in Sec. IV-A. Given the input discussed in Sec. III, a sequential task plan and the corresponding task graph, as shown in Fig. 2 (a), are generated from the assembly using ILP. In the next stage, explained in Sec. IV-B, the motion plan for each task is generated sequentially with a single-robot motion planner (see Fig. 2 (b)). Sec. IV-C presents the last and most crucial stage, which converts the sequential plan to a multi-modal TPG (e.g. Fig. 2 (c)) that can be executed safely, asynchronously, and efficiently in real robots.
For notation, we use $i\in[1,N]$ to index robots, $j\in[1,N_{a}]$ to index assembly steps in $A$, $m$ to index tasks for each robot, and $k\in[1,N_{b}]$ to index moveable objects in $\mathcal{B}$.
@@ IV-A Task Planning
Given the assembly sequence $A$, task planning aims to construct a sequential task plan that includes robot assignment, robot target pose assignment, and object assignment. Specifically, for each step $a_{j}$ in the assembly sequence, the task planner must assign one responsible robot, a support robot if required, and an object $b_{k}$ of the correct type. In addition, the task planner must select the feasible grasp pose and a feasible support pose if required to perform the necessary manipulation skills for each $a_{j}$. Fig. 3 provides an overview of all the decisions made. These decisions directly affect the feasibility of motion planning, as task planning determines whether there are collision-free and deadlock-free paths.
@@
Task Definition We define a task $\mathcal{T}^{i}$ as a piece of work that requires a robot $i$ to either transit to a goal pose or to perform some manipulation skill $s$ to achieve a goal constraint. For example, a task may involve a robot arm picking up an object, supporting a structure, receiving an object handed from another robot, retracting to its HOME pose, etc. In the context of LEGO assembly, each assembly task $a_{j}$ is broken down into a sequence of tasks. As shown in Fig. 1 (a) and (b), a robot must first transit to the initial location of an object, pick the object, transit the object to the location for assembly, place down the object on the target structure, and finally transit back to the robot’s HOME pose. For an assembly step that requires collaborative assembly, the support robot is also assigned a sequence of tasks, such as transiting to the structure and supporting it in Fig. 1 (c) and (e). If a reorientation is required, as illustrated in Fig. 1 (d), the support robot is assigned a transit task and a pick task to collect the object, and a handover task to the other robot so that the object can be placed up to the structure.
@@
Task Graph A task graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a direct acyclic graph that represents the ordered set of tasks for all robots to complete the final assembly. The task graph also implicitly represents the movement of manipulated objects, the evolving environment and planning scene, as well as the kinematics switches of the robots. A node is either a task node or an object node. A task node represents a task $\mathcal{T}^{i}_{m}\in\mathcal{V}$ for robot $i$, where $m$ is the index. Each task node also contains a robot target pose for this task. An object node represents an object $b_{k}$ and its pose. An edge from one task node to another in the graph, $\mathcal{T}^{i}_{m}\rightarrow\mathcal{T}^{i^{\prime}}_{m^{\prime}}$, represents a precedence constraint that states a dependency where $\mathcal{T}^{i}_{m}$ must finish before $\mathcal{T}^{i^{\prime}}_{m^{\prime}}$ can start. An edge from an object node to a task node, $b_{k}\rightarrow\mathcal{T}^{i}_{m}$, means that the object $b_{k}$ is kinematically attached to the robot $i$ at the beginning of this task. In contrast, an edge from a task node to an object node, $\mathcal{T}^{i}_{m}\rightarrow b_{k}$, represents that an object would be detached from robot $i$ after this task ends. Each robot $i$ has $M^{i}$ tasks. A task graph itself does not limit whether the tasks must be executed sequentially, synchronously, or asynchronously.
@@
Approach The main idea of our approach is to find a sequential turn-based task plan according to the assembly sequence $A$, which is itself sequential. Only one robot is actively moving or executing skills at any time, while the other robot waits. Each robot returns to its HOME pose at the end of completing an assembly step.
@@
Algorithmically, an ILP jointly optimizes robot assignment, object assignment, and target robot poses. A set of binary decision variables assigns a robot $i$ to an assembly step $a_{j}$ using the object $b_{k}$ and a feasible grasp pose. Another set of binary decision variables denotes the assignment of the support robot and support poses. We precompute feasible robot poses for grasping all available objects at initial and assembled positions, and support poses if necessary. The cost of assigning a robot $i$, an object $b_{k}$, and a corresponding grasp pose to an assembly step $a_{j}$ is estimated by the sum of the transit distances necessary for this step. The ILP finds the best set of assignments that minimizes the sum of costs to complete the assembly and an auxiliary term for load balancing while ensuring that the object type and support robot requirements are met. More details of the ILP formulation are discussed in Sec. -B.
Combined with the precomputed robot poses, the optimized assignment gives a complete set of robot, object, and grasp pose assignments in each step. We then construct a corresponding sequential task plan and task graph (e.g., Fig. 2 (a)) with all inter-robot task dependencies and object relationships. A sequence of tasks for the assigned robot is added for each assembly step, and two object nodes are added to the task graph and connected to the pick task and place task nodes to indicate their attachment and placement, respectively. If a collaborative assembly is required, a sequence of support tasks is added to the support robot to be completed first. Inter-robot task dependencies are added to the task graph to constrain that the support task must precede any place task, and the following task after support can only start after the place task finishes. If two consecutive assembly steps $a_{j}$ and $a_{j+1}$ are assigned to different robots, a task dependency is added to ensure that the place task for $a_{j+1}$ must wait for the place task for $a_{j}$.
Compared to other MR-TAMP and assembly methods such as [13, 18, 42, 5], APEX-MR first generates a sequential plan. This has two advantages: (1) It is easy to reason about inter-robot collision for scheduling tasks and avoids expensive feasibility checks needed for parallel task execution, thereby making the ILP easier to solve; (2) The complexity of motion planning is significantly reduced since each task becomes a single-robot planning problem, which avoids solving a challenging and time-consuming multi-robot arm motion planning problem. It is worth noting that, while the sequential plan might seem inefficient for a multi-robot system, our TPG execution framework will post-process this plan to enable efficient parallel execution.
@@ IV-B Motion Planning
Once the sequential task plan for each robot is determined, motion planning becomes straightforward. As illustrated in Fig. 2 (b), APEX-MR iterates over each task and uses a single-robot RRT-Connect algorithm to plan the path for this task when other robots are waiting. This is feasible because all other robots would be at a nonblocking stationary pose, i.e., HOME [^1] We also assume that any robot waiting at a support pose is nonblocking.. A reference path is also generated for tasks executed by specific manipulation skills based on the grasp/support pose.
The motion planner generates a planned path for every robot and every task. Each path $\tau^{i}_{m}$ for task $\mathcal{T}_{m}^{i}$ of robot $i$ is represented with a sequence of uniformly timestamped poses, $\{(C_{n}^{i},t^{i}_{n})\},n\in[N^{i}_{start,m},N^{i}_{end,m}]$ with step size $\Delta t$. $N^{i}_{start,m}$ and $N^{i}_{end,m}$ are the first and last index of the path $\tau^{i}_{m}$, $N^{i}_{start,1}=1$, and $N^{i}_{start,m+1}=N^{i}_{end,m}+1$ for $m=1,\dots M-1$. For each task, the planner first generates a sequence of poses ${C^{i}_{n}}$ for the robot $i$, then determines the corresponding timesteps. To ensure that each robot takes turns to complete its task according to the task sequences, the initial timestep of each task is equal to the last timestep of the previous task in the sequential plan. Then, the timestep $t_{n}^{i}$ for each pose of the same task is set based on the maximum robot velocity, i.e., $t^{i}_{n}=t^{i}_{n-1}+d(C^{i}_{n},C^{i}_{n-1})/v_{max}$. Since RRT-Connect may produce jerky and long paths, we use a randomized shortcutting algorithm to smooth suboptimal paths. The output of motion planning should be a sequence of paths, one for each task, i.e., $\tau^{i}_{m},\;\forall i\in\{1,\dots,N\},\;\forall m\in\{1,\dots,M^{i}\}$.
@@ IV-C Asynchronous Execution
TPG Definition APEX-MR uses a multi-modal temporal plan graph (TPG) to represent an execution schedule for the team of robot arms. As illustrated in Fig. 2 (c), multi-modal TPG is a directed acyclic graph $G=(V,E)$ with two types of nodes. A pose node $v_{n}^{i}$ corresponds to a configuration $C_{n}^{i}$ on the path of a transit task for robot $i$. A skill node $\tilde{v}^{i}_{n}$ represents a manipulation skill (e.g., pick, handover, support, etc.) that will be executed by some robot controller or policy. The skill node also contains a reference path generated in motion planning. A type-1 edge $(v_{n}^{i}\rightarrow v_{n+1}^{i})$ [^2] Type-1 edge connects skill nodes too, i.e., $(\tilde{v}_{n}^{i}\rightarrow v_{n+1}^{i})$ or $(v_{n}^{i}\rightarrow\tilde{v}_{n+1}^{i})$ connects two nodes from the same robot $i$ and indicates the order between two nodes. A type-2 edge $(v_{n}^{i}\rightarrow v_{n^{\prime}}^{i^{\prime}})$ represents an inter-robot precedence order that constrains robot $i^{\prime}$ to wait for robot $i$ to reach $v_{n}^{i}$ (i.e., reach pose $C_{n}^{i}$ or complete the manipulation skill) before moving to $v_{n^{\prime}}^{i^{\prime}}$. A type-2 edge can be added for both task dependencies and motion dependencies. For every pair of TPG nodes that could lead to collisions, this is a motion dependency, and a type-2 edge is required. TPG ensures a deadlock-free execution schedule if and only if there are no cycles in the TPG. In contrast to the TPG defined in [19], a multi-modal TPG combines a TPG with a task graph and assigns a corresponding task $\mathcal{T}_{m}^{i}$ to each node $v_{n}^{i}$. Since there are kinematics switches and changes to the planning scene, each node also contains the robot kinematics (i.e., attached object), which will be important for the TPG construction process.
@@
Input: A task graph $\mathcal{G}$and all robot paths $\tau^{i}_{m}$robot $i=1,\dots,N$task $\mathcal{T}_{m}^{i}=\mathcal{T}^{i}_{1},\dots,\mathcal{T}_{M^{i}}^{i}$task $\mathcal{T}_{m}^{i}$is a transit to a goal pose
1:Add a sequence of nodes $\{v^{i}_{n}\}$ from $\tau_{m}^{i}$
@@
9:Simplify TPG with transitive reduction
@@
Building a multi-modal TPG The process of constructing a TPG from a multi-robot motion plan can be interpreted as converting from the sequential robot paths to temporally dependent robot schedules. On a high level, inter-robot task dependencies are copied as type-2 edges for TPG, and type-2 edges for motion dependencies are constructed by scanning for collisions between all pairs of TPG nodes. A key benefit of this partial-order representation is that TPG does not specify a fixed time between two consecutive nodes, and thus allows execution delays. We outline the detailed procedure below.
@@
The construction process begins with creating the nodes and type-1 edges. For each transit task $\mathcal{T}_{m}^{i}$, a pose node $v_{n}^{i}$ is constructed for each configuration $C_{n}^{i}$ of the path $\tau_{m}^{i}$. A skill node $\tilde{v}_{n}^{i}$ is created for every manipulation task. The timestamp in the input path $t_{n}^{i}$ for each node $v_{n}^{i}$ is also recorded. Each node $v_{n}^{i}$ is then connected to its successor $v_{n+1}^{i}$ by a type-1 edge, and the end node of a task $v_{end,m}^{i}$ is connected to the start node of the next task $v_{start,m+1}^{i}$. When a robot waits at a node $v_{m}^{i}$, that is $C_{m}^{i}=C_{m-1}^{i}$, the node $v_{m}^{i}$ is removed because removing the wait action does not change the robot path or any temporal dependency in the TPG.
@@
Next, all the type-2 edges are identified in the TPG. First, all inter-robot task dependencies $(\mathcal{T}_{m}^{i}\rightarrow\mathcal{T}_{m^{\prime}}^{i^{\prime}})\in\mathcal{E}$ from the task graph $\mathcal{G}$ are added as type-2 edges to the TPG. Specifically, for every edge in the task graph $\mathcal{T}_{m}^{i}\rightarrow\mathcal{T}_{m^{\prime}}^{i^{\prime}}$, a type-2 edge is added from the beginning of task $\mathcal{T}_{m+1}^{i}$, $v_{{start,m+1}}^{i}$, to the beginning of task $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$, $v_{{start,{m^{\prime}}}}^{i^{\prime}}$. This type-2 edge ensures that robot $i$ only starts $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$ after robot $i^{\prime}$ finishes task $\mathcal{T}_{m}^{i}$.
@@
Then, motion-level dependencies are identified by iterating over all pairs of nodes in the TPG with a double loop iterating over $(i,i^{\prime})$. Robot kinematics can be incrementally changed if node $v_{n}^{i}$ is the start of a new task that attaches or detaches an object. For each node $v_{n}^{i}$, it is checked against all other nodes $v_{n^{\prime}}^{i^{\prime}}$ from a different robot ($i\neq i^{\prime}$) and has an earlier timestamp $t^{i^{\prime}}_{n^{\prime}}\leq t_{n}^{i}$. Collision checking for pose nodes means that the robot links and attached objects at the corresponding poses $C_{n}^{i}$ and $C_{n^{\prime}}^{i^{\prime}}$ are checked. For skill nodes, all robot poses on the reference path have to be collision-free simultaneously. If there is any collision between $v_{n^{\prime}}^{i^{\prime}}$ and $v_{n}^{i}$, a type-2 edge is added from the earlier node’s successor, $v_{n^{\prime}+1}^{i^{\prime}}$, to the later node, $v_{n}^{i}$. Setting the direction of type-2 edge based on timestamps in the sequential motion plan ensures that the multi-modal TPG has no cycles and thus is deadlock-free. This way, if $v_{n^{\prime}}^{i^{\prime}}$ has a smaller timestamp than $v_{n}^{i}$ in the input sequential plan, i.e., $t_{n^{\prime}}^{i^{\prime}}<t_{n}^{i}$, $v_{n^{\prime}}^{i^{\prime}}$ must still be executed before $v_{n}^{i}$ can start, avoiding any potential collisions. When iterating $n^{\prime}$, it is unnecessary to check collisions for any node $v_{n^{\prime}}^{i^{\prime}}$ that has a larger timestamp than $v_{n}^{i}$, i.e., $t_{n^{\prime}}^{i^{\prime}}\geq t_{n}^{i}$, because those would be checked when the iterated robot pair ($i$, $i^{\prime}$) are swapped. This ensures the eventual multi-modal TPG is cycle-free and hence deadlock-free, since the Also, if the current $v_{n^{\prime}}^{i^{\prime}}$ is a predecessor of $v_{n}^{i}$ in the graph, $v_{n}^{i}$ already waits for $v_{n^{\prime}}^{i^{\prime}}$ and avoid collisions, so it becomes unnecessary to check them again.
@@
Figure 4: Illustration of the execution setup. TPG maintains and controls the execution schedule of all robots on a central server, gradually sending new path segments that can be safely executed. Each robot maintains a controller-sensing loop independently while updating its state with the central server.
Further Optimization An optional step to further reduce the execution makespan and smooth the path is to skip the intermediate transition to HOME after every assembly step. We use the following shortcutting algorithm, similar to the strategy implemented in [41], to achieve that while maintaining a collision- and deadlock-free plan.
The anytime algorithm works by randomly sampling two nodes of the TPG ($v_{n}^{i},v_{n^{\prime}}^{i}$) from the same transit task $\mathcal{T}_{m}^{i}$, checking whether connecting them in a shortcut is feasible. Consecutive transit tasks passing through the robot’s HOME pose are merged as a single task. This allows the HOME pose to be skipped. A shortcut path directly interpolates between $C_{n}^{i}$ and $C_{n^{\prime}}^{i}$ to generate a sequence of poses with the same step size $\Delta t$. The shortcut must be collision-checked against any independent nodes (i.e., nodes that are not predecessors to $v_{n}^{i}$ or successors to $v_{n^{\prime}}^{i}$ in the TPG). On a multi-modal TPG, the collision checking must include any attached objects to the robot, as well as any independent object nodes on a task graph (i.e., object nodes that are not predecessors or successors of the current task). Once a valid shortcut is found, the original nodes between $v_{n}^{i}$ and $v_{n^{\prime}}^{i}$ are replaced by new nodes corresponding to the shortcut. If adding a valid shortcut from $v_{n}^{i}$ to $v_{n^{\prime}}^{i}$ skips any outgoing edges between these two nodes, the start nodes of these outgoing edges are moved to $v_{n}^{i}$. If any incoming edges are skipped, then the end nodes of these incoming edges are moved to $v_{n^{\prime}}^{i}$. These two steps ensure that any dependencies before adding a shortcut still exist after, and the TPG remains collision-free. Since each shortcut is collision-free, no new dependencies in the TPG are introduced, and TPG remains deadlock-free. The shortcutting algorithm keeps identifying valid shortcuts until a user-defined time limit is reached, removing redundant transitions to HOME poses in the process.